1)Ivana invested her savings of $4800. SHe invested part in a mutual fund, which paid her 9% interest per year, and the rest went into a GIC, which paid her 10% per year. After one year, the interest she earned from the mutual fund was $43 less than the interest she earned from the GIC. How much did Hailey invest into each?
Answer by mananth(16949)

Let her invest $ x in Mutual funds
& $y in GIC
x+y= 4800
.
Interest on mutual fund = 9%
Interest accrued = 0.09x
.
Interest on GIC is 10%
Interest accrued = 0.1y
.
0.09x=0.1y -43
0.09x -0.1y =-43 ( multiply by 100)
.
9x -10y=-4300
10x+10y+9x-10y = 48000-4300
19x=43700
x= $2300 the amount she invested in Mutual fund.
y=$2500 in GIC